# Optimizing Information Freshness of IEEE 802.11ax Uplink OFDMA-Based Random Access

### v1

@@ V-A Theoretical Analysis
$\begin{split}\overline{\Delta}&=\frac{\frac{1}{6}\alpha_{0}(\alpha_{0}+1)(2\alpha_{0}+1)L+(\alpha_{0}+1)^{2}\beta_{0}+1}{(\alpha_{0}+1)\alpha_{0}+2(\alpha_{0}+1)\beta_{0}+2}\\ &+\left(\left(1-\frac{1}{L\mathbb{E}[U_{0}]}\right)^{1-N}-1\right)\mathbb{E}[U_{0}]+\frac{1}{2},\end{split}$ (34)
@@
$\begin{split}\overline{\Delta}_{LB}&=\left(\left(1-\frac{1}{L\mathbb{E}[U_{0}]}\right)^{1-N}-\frac{1}{2}\right)\mathbb{E}[U_{0}]+\frac{1}{2}.\end{split}$ (35)
@@
The expression of $\overline{\Delta}$ given in Theorem 4 is a piecewise multivariate multi-order function of $\alpha_{0}$ and $\beta_{0}$, which is significantly challenging to investigate thoroughly. Since $\overline{\Delta}_{LB}$ is more analytically feasible than $\overline{\Delta}$, we are motivated to minimize this lower bound of $\overline{\Delta}$ to guide the UORA parameter optimization. Then, we find that the complexity of the expression structure of $\overline{\Delta}_{LB}$ renders it intractable to comprehend, thereby limiting the theoretical insights it offers. Hence, we consider $\overline{\Delta}_{LB}$ in a special case when $W_{0}-1$ is divisible by $L$, i.e., $W_{0}-1=L\alpha_{0}$. Then, (35) can be rewritten and approximated as
@@
where $(a)$ is obtained by applying the approximation $(1-x)^{n}\approx\exp(-nx)$ for $x\in(0,1)$ and large $n\in\mathbb{N}$ to the term $q=(1-\frac{\rho}{L})^{N-1}$. Compared with (35), $\tilde{\Delta}_{LB}$ has fewer independent variables and degrees, and thus is more tractable for analysis.
@@ V-B Optimization Algorithm Design
Recall that we aim to minimize the lower bound of $\overline{\Delta}$ to guide the UORA parameter optimization. This can be achieved by using the behavior of $\overline{\Delta}_{LB}(W)$, defined as (35) by substituting $\alpha_{0}=\lfloor(W-1)/L\rfloor$ and $\beta_{0}=W-1-\alpha_{0}L$. Note that we only obtain the behavior of an approximation of $\overline{\Delta}_{LB}(W)$ in $(0,+\infty)$ when $W-1$ is divisible by $L$, namely $\tilde{\Delta}_{LB}(W)$. However, it is not immediately evident that $\tilde{\Delta}_{LB}(W)$ exhibits similar behavior to $\overline{\Delta}_{LB}(W)$. To proceed, we define function $\hat{\Delta}_{LB}(W)$ as the left hand side of (36) by replacing $W_{0}$ with $W$ and define
@@ VII-D Proof of Theorem 4
Substituting (58) and (59) into (57), we obtain (34) and (35).

### v2

@@ V-A Theoretical Analysis
$\overline{\Delta}$ $=\frac{\mathbb{E}[U_{0}^{2}]}{2\mathbb{E}[U_{0}]}+\frac{1-q}{q}\mathbb{E}[U_{0}]+\frac{1}{2}$ (34a)
$=\frac{\frac{1}{6}\alpha_{0}(\alpha_{0}+1)(2\alpha_{0}+1)L+(\alpha_{0}+1)^{2}\beta_{0}+1}{(\alpha_{0}+1)\alpha_{0}+2(\alpha_{0}+1)\beta_{0}+2}$ (34b)
$+\left(\left(1-\frac{1}{L\mathbb{E}[U_{0}]}\right)^{1-N}-1\right)\mathbb{E}[U_{0}]+\frac{1}{2},$
@@
$\overline{\Delta}_{LB}$ $=\frac{\mathbb{E}[U_{0}]^{2}}{2\mathbb{E}[U_{0}]}+\frac{1-q}{q}\mathbb{E}[U_{0}]+\frac{1}{2}$ (35a)
$=\left(\left(1-\frac{1}{L\mathbb{E}[U_{0}]}\right)^{1-N}-\frac{1}{2}\right)\mathbb{E}[U_{0}]+\frac{1}{2}.$ (35b)
@@
The expression of $\overline{\Delta}$ given in Theorem 4 is a piecewise multivariate multi-order function of $\alpha_{0}$ and $\beta_{0}$, which is significantly challenging to investigate thoroughly. To proceed, we consider its lower bound $\overline{\Delta}_{LB}$. Clearly, the only difference between the expressions of $\overline{\Delta}$ and $\overline{\Delta}_{LB}$, given in (34a) and (35a), is the numerator of the first term. For (34a), the numerator of the first term is $\mathbb{E}[U_{0}^{2}]$, and for (35a), the numerator is $\mathbb{E}[U_{0}]^{2}$. Our numerical results show that the ratio $(\mathbb{E}[U_{0}])^{2}/\mathbb{E}[U_{0}^{2}]$ is greater than $73\%$ under all feasible UORA configurations and this ratio increases as $L$ increases. We thus conjecture that the deviation between $\overline{\Delta}$ and $\overline{\Delta}_{LB}$, defined as $(\overline{\Delta}-\overline{\Delta}_{LB})/\overline{\Delta}$, could be insignificant. We have validated this conjecture through numerical results, which show that the deviation remains below $5\%$ when $N$ and $L$ go large. This implies that $\overline{\Delta}$ and $\overline{\Delta}_{LB}$ have similar behaviors. Built upon this observation, we are motivated to minimize the lower bound $\overline{\Delta}_{LB}$ to guide the UORA parameter optimization. Then, we find that the complexity of the expression structure of $\overline{\Delta}_{LB}$ renders it intractable to comprehend, thereby limiting the theoretical insights it offers. Hence, we consider $\overline{\Delta}_{LB}$ in a special case when $W_{0}-1$ is divisible by $L$, i.e., $W_{0}-1=L\alpha_{0}$. Then, (35b) can be rewritten and approximated as
@@
where $(a)$ is obtained by applying the approximation $(1-x)^{n}\approx\exp(-nx)$ for $x\in(0,1)$ and large $n\in\mathbb{N}$ to the term $q=(1-\frac{\rho}{L})^{N-1}$. Compared with (35b), $\tilde{\Delta}_{LB}$ has fewer independent variables and degrees, and thus is more tractable for analysis.
@@ V-B Optimization Algorithm Design
Recall that we aim to minimize the lower bound of $\overline{\Delta}$ to guide the UORA parameter optimization. This can be achieved by using the behavior of $\overline{\Delta}_{LB}(W)$, defined as (35b) by substituting $\alpha_{0}=\lfloor(W-1)/L\rfloor$ and $\beta_{0}=W-1-\alpha_{0}L$. Note that we only obtain the behavior of an approximation of $\overline{\Delta}_{LB}(W)$ in $(0,+\infty)$ when $W-1$ is divisible by $L$, namely $\tilde{\Delta}_{LB}(W)$. However, it is not immediately evident that $\tilde{\Delta}_{LB}(W)$ exhibits similar behavior to $\overline{\Delta}_{LB}(W)$. To proceed, we define function $\hat{\Delta}_{LB}(W)$ as the left hand side of (36) by replacing $W_{0}$ with $W$ and define
@@ VII-D Proof of Theorem 4
Substituting (58) and (59) into (57), we obtain (34b) and (35b).